# Markov Perfect Equilibria in

### v1

@@ 7 Continuous-time limit
Step 1. We claim that $(v^{(K)},\mathfrak{a}^{(K)}/h^{(K)})_{K\geq K_{0}}$ is a uniformly bounded, equicontinuous family of continuous functions, seen as functions from $[0,T]$ to $[0,\infty)^{\mathscr{X}_{N}}\times(\mathbb{R}^{d})^{\mathscr{X}_{N}}$.
Uniform boundedness. By Assumption 2.5, $v^{(K)}$ are uniformly bounded, say by some $c_{0}<\infty$. We set
$\tilde{H}(x,z,v,\tilde{a}):=\ell(x,z,a/h)h+\sum_{y\neq x}a(y)\Delta_{x}v(y)+v(x),\quad(x,z,v,a)\in\mathscr{X}_{N}\times\mathbb{R}^{d}\times\mathbb{R}^{d-1},$
and note that $\tilde{H}(x,z,v,a_{-x})=H(x,z,v,a)$ for all $(x,z,v,a)\in\mathscr{X}_{N}\times\mathbb{R}^{d-1}\times\Sigma^{d-1}$. We note that $\alpha^{(K)}_{-x}(s^{(K)}_{k},x,z)$ is given as the minimizer of the Hamiltonian
$\tilde{A}(h^{(K)})\ni\tilde{a}_{-x}\mapsto\tilde{H}(x,z,E(x,z,v^{(K)}(s^{(K)}_{k+1}),\alpha^{(K)}(s^{(K)}_{k})),\tilde{a}_{-x}),$
where $\tilde{A}(h^{(K)})$ is the set of subprobability vectors $\tilde{a}_{-x}\in[0,1]^{d-1}$ satisfying $\sum_{y\neq x}\tilde{a}_{x}\leq 1$ and for which there exists $a\in A(h^{(K)})$ with $a_{-x}=\tilde{a}_{-x}$. Let $\eta^{(K)}$ be defined as the minimizer of $\tilde{A}(h^{(K)})\ni\tilde{a}_{-x}\mapsto\tilde{H}(x,z,0,\tilde{a}_{-x})$. By Lemma 3.3,
$\frac{1}{h^{(K)}}|\alpha^{(K)}_{-x}(s^{(K)}_{k},x,z)-\eta^{(K)}(x,z)|_{1}\leq\frac{1}{\gamma}\,\sum_{y\in S(x)\setminus\{x\}}\mathbb{E}|\Delta_{x}v^{(K)}(s^{(K)}_{k+1},x,Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k})))|\leq\frac{2(m-1)c_{0}}{\gamma}.$
By definition of $\eta^{(K)}$ as the minimizer of $\tilde{H}(x,z,0,\cdot)$ over $\tilde{A}(h^{(K)})$,
$\ell(x,z,\frac{\eta^{(K)}(x,z)}{h^{(K)}})\leq\ell(x,z,\sigma^{2}).$
Since $\ell(x,z,\cdot)$ is strongly convex, it is coercive, and hence $(|\eta^{(K)}(x,z)|_{1}/h^{(K)})_{K\geq 1}$ is bounded by some $\tilde{c}_{0}<\infty$. Putting these two estimates together, we arrive at
$\sup_{K\geq K_{0}}\max_{0\leq k<K}\max_{(x,z)\in\mathscr{X}_{N}}\frac{1}{h^{(K)}}|\alpha_{-x}^{(K)}(s^{(K)}_{k},x,z)|_{1}\leq\frac{2(m-1)c_{0}}{\gamma}+\tilde{c}_{0}=:c_{1}<\infty.$
Using the relation
$|\mathfrak{a}^{(K)}(s^{(K)}_{k},x,z)|_{1}=2|\alpha_{-x}^{(K)}(s^{(K)}_{k},x,z)|_{1},$
this implies the claim. Note that $(|\alpha^{(K)}(t,x,z)|_{1}/h^{(K)})_{K\geq K_{0}}$ is not a bounded sequence since $\alpha^{(K)}_{x}(t,x,z)\to 1$ as $K\to\infty$.
Equicontinuity. Let us first observe that by the same reasoning as in Lemma 7.1, the uniform boundedness of $(\mathfrak{a}^{(K)}/h^{(K)})_{K\geq K_{0}}$ implies that
@@
$|v(s^{(K)}_{k+1},x,z)-v(s^{(K)}_{k},x,z)|$
$\quad\leq|\mathbb{E}^{\alpha^{(K)}\otimes\alpha^{(K)}}[v(s^{(K)}_{k+1},X_{s^{(K)}_{k+1}},Z_{s^{(K)}_{k+1}})]-v(s^{(K)}_{k},x,z)|+|\mathbb{E}^{\alpha^{(K)}\otimes\alpha^{(K)}}[v(s^{(K)}_{k+1},X_{s^{(K)}_{k+1}},Z_{s^{(K)}_{k+1}})]-v(s^{(K)}_{k+1},x,z)|$
$\quad=|\mathbb{E}^{\alpha^{(K)}\otimes\alpha^{(K)}}[v(s^{(K)}_{k+1},X_{s^{(K)}_{k+1}},Z_{s^{(K)}_{k+1}})]-v(s^{(K)}_{k},x,z)|$
$\quad+\sum_{(x^{\prime},z^{\prime})\neq(x,z)}\mathbb{P}^{\alpha^{(K)}\otimes\alpha^{(K)}}[(X_{s^{(K)}_{k+1}},Z_{s^{(K)}_{k+1}})=(x^{\prime},z^{\prime})\,|\,(X_{s^{(K)}_{k}},Z_{s^{(K)}_{k}})=(x,z)]\,|v(s^{(K)}_{k+1},x^{\prime},z^{\prime})-v(s^{(K)}_{k+1},x,z)|$
$\quad\leq|\ell(x,z,\alpha^{(K)}(s_{k}^{(K)},x,z)/h^{(K)})|h^{(K)}+2c_{0}c_{2}\,h^{(K)}\leq(\sup_{[0,c_{1}]^{d-1}}|\ell(x,z,\cdot)|+2c_{0}c_{2})\,h^{(K)}=:c_{3}h^{(K)}.$
This implies equicontinuity of $(v^{(K)})_{K\geq K_{0}}$. We turn to show the equicontinuity of $(\mathfrak{a}^{(K)})_{K\geq K_{0}}$. As in the proof in Proposition 3.4, by Lemma 3.3,
$|\alpha^{(K)}_{-x}(s^{(K)}_{k},x,z)$ $\,-\alpha^{(K)}_{-x}(s^{(K)}_{k+1},x,z)|_{1}$
$\leq\frac{h^{(K)}}{\gamma}\sum_{y\in S(x)\setminus\{x\}}\mathbb{E}|\Delta_{x}v(s^{(K)}_{k+1},y,Z^{\prime}(x,z,\alpha^{(K)}(s_{k})))-\Delta_{x}v(s^{(K)}_{k+2},y,Z^{\prime}(x,z,\alpha^{(K)}(s_{k+1})))|$
$\leq\frac{4(m-1)L^{(K)}_{k}h^{(K)}}{\gamma}\max_{(x,z)\in\mathscr{X}_{N}}|\alpha^{(K)}_{-x}(s^{(K)}_{k},x,z)-\alpha^{(K)}_{-x}(s^{(K)}_{k+1},x,z)|_{1}+\frac{c_{3}(m-1)}{\gamma}(h^{(K)})^{2},$
@@
$\bigl{(}1-\frac{4(m-1)L^{(K)}_{k}h^{(K)}}{\gamma}\bigr{)}\max_{(x,z)\in\mathscr{X}_{N}}|\alpha^{(K)}_{-x}(s^{(K)}_{k},x,z)-\alpha^{(K)}_{-x}(s^{(K)}_{k+1},x,z)|_{1}\leq\frac{c_{3}(m-1)}{\gamma}(h^{(K)})^{2}$
@@
Step 2. By the Arzelà-Ascoli theorem, there exists continuous $(\tilde{v}^{(\infty)},\tilde{\alpha}^{(\infty)})$ such that, along a subsequence, $(v^{(K)},\mathfrak{a}^{(K)}/h_{K})\to(\tilde{v}^{(\infty)},\tilde{\alpha}^{(\infty)})$ uniformly on $[0,T]$ as $K\to\infty$. We replace the original sequence with this subsequence.
Step 3. For $(x,z)\in\mathscr{X}_{N}$ define
$\varphi^{(K)}_{k}(x,z)$ $:=\tilde{H}(x,z,E(x,z,v^{(K)}(s^{(K)}_{k+1}),\alpha^{(K)}(s^{(K)}_{k})),\mathfrak{a}^{(K)}(s^{(K)}_{k},x,z))-v^{(K)}(s^{(K)}_{k+1},x,z),\quad 0\leq k<K$
$\varphi^{(\infty)}(s,x,z)$ $:={\mathcal{H}}(x,z,\Delta_{x}\tilde{v}^{(\infty)}(s,\cdot,z),\tilde{\alpha}^{(\infty)}(s,x,z))+\mathscr{L}^{\tilde{\alpha}^{(\infty)}(s)}\tilde{v}^{(\infty)}(s,\cdot,\cdot)(x,z)),\quad s\in[0,T].$
We claim that
$\lim_{K\to\infty}\ \max_{0\leq k<K}\ \Bigl{|}\frac{1}{h^{(K)}}\varphi^{(K)}_{k}(x,z)-\varphi^{(\infty)}(s^{(K)}_{k},x,z)\Bigr{|}=0.$
Indeed, write $\varphi^{(K)}_{k}:=\ell(x,z,\mathfrak{a}^{(K)}(s^{(K)}_{k},x,z)/h^{(K)})h^{(K)}+\mathrm{(I)}_{k}+\mathrm{(II)}_{k}$ where
$\mathrm{(I)}_{k}$ $:=\sum_{y\in{\mathcal{X}}\setminus\{x\}}\mathfrak{a}^{(K)}_{y}(s^{(K)}_{k},x,z)\,\mathbb{E}[v^{(K)}(s^{(K)}_{k+1},y,Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k})))-v^{(K)}(s^{(K)}_{k+1},x,Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k})))],$
@@
Using the fact that $\lim_{K\to\infty}Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k}))=z$ and the uniform convergence of $(v^{(K)},\mathfrak{a}^{(K)}/h^{(K)})$ to $(\tilde{v}^{(\infty)},\tilde{\alpha}^{(\infty)})$,
$\lim_{K\to\infty}\ \max_{0\leq k<K}\ \Bigl{|}\frac{1}{h^{(K)}}\mathrm{(I)}_{k}-\sum_{y\in{\mathcal{X}}\setminus\{x\}}\tilde{\alpha}^{(\infty)}_{y}(s^{(K)}_{k},x,z)\,(\tilde{v}^{(\infty)}(s^{(K)}_{k},y,z)-\tilde{v}^{(\infty)}(s^{(K)}_{k},x,z))\Bigr{|}=0.$
@@
$\mathrm{(II)}_{k}=\sum_{z^{\prime}\in\Sigma^{d-1}_{N}\setminus\{z\}}\mathbb{P}[Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k}))=z^{\prime}]\,(v^{(K)}(s^{(K)}_{k+1},x,z^{\prime})-v^{(K)}(s^{(K)}_{k+1},x,z)),$
and we use Lemma 7.1 and the uniform convergence of $(v^{(K)},\mathfrak{a}^{(K)}/h^{(K)})$ to $(\tilde{v}^{(\infty)},\tilde{\alpha}^{(\infty)})$ to see
$\lim_{K\to\infty}\ \max_{0\leq k<K}\ \Bigl{|}\frac{1}{h^{(K)}}\mathrm{(II)}_{k}-\sum_{w\neq y}Nz_{y}\,\tilde{\alpha}^{(\infty)}_{w}(s^{(K)}_{k},y,z+e^{N}_{xy})\,(\tilde{v}^{(\infty)}(s^{(K)}_{k},x,z+e^{N}_{wy})-\tilde{v}^{(\infty)}(s^{(K)}_{k},x,z))\Bigr{|}=0.$
Step 3. Note that for every $0\leq k<K$ we have
$-(v^{(K)}(s^{(K)}_{k+1},x,z)-v^{(K)}(s^{(K)}_{k},x,z))=\varphi^{(K)}_{k}(x,z),\quad(x,z)\in\mathscr{X}_{N}.$ (7.1)
Let $t\in(0,T)$, choose a sequence $t^{(K)}\in\{0,T/K,\ldots,T\}$ converging to $t$ and let $T^{(K)}=t^{(K)}K$. Summing over $10\leq k\leq T^{(K)}-$ 1 in (7.1), we obtain $-(v^{(K)}(t^{(K)},x,z)-v^{(K)}(0,x,z))=\sum_{k=0}^{T^{(K)}-1}\varphi^{(K)}_{k}(x,z)$. Now
@@
$+\Bigl{|}\sum_{k=0}^{T^{(K)}-1}\varphi^{(\infty)}(s^{(K)}_{k},x,z)h_{K}-\int_{0}^{t}\varphi^{(\infty)}(s,x,z)\,\mathrm{d}s\Bigr{|}$
The first term converges to zero by Step 2. while the second term converges to zero by classical arguments. This proves that $\tilde{v}^{(\infty)}(t,x,z)-\tilde{v}^{(\infty)}(0,x,z)=-\int_{0}^{t}\varphi^{(\infty)}(s,x,z)\,\mathrm{d}s$. Hence $\tilde{v}^{(\infty)}(\cdot,x,z)$ is continuously differentiable on $[0,T]$ with derivative $-\varphi^{(\infty)}(\cdot,x,z)$.
Step 4. We claim that
$(\tilde{\alpha}^{(\infty)}(t,x,z))_{y\neq x}=\underset{a\in[\sigma^{2},\infty)^{d-1}}{\mathrm{arg\,min}}\ {\mathcal{H}}(x,z,\Delta_{x}\tilde{v}^{(\infty)}(t,\cdot,z),a),\quad(t,x,z)\in[0,T]\times\mathscr{X}_{N}.$
Then, for every $0\leq k<K$ and $(x,z,a)\in\mathscr{X}_{N}\times\tilde{A}(h^{(K)})$ we have,
$\tilde{H}(x,z,E(x,z,v^{(K)}(t^{(K)}+h),\alpha^{(K)}(t^{(K)})),\mathfrak{a}^{(K)}(t^{(K)},x,z))\leq\tilde{H}(x,z,E(x,z,v^{(K)}(t^{(K)}+h),\alpha^{(K)}(t^{(K)})),a).$
Adding $\mathbb{E}[v(t^{(K)}+h,x,Z^{\prime}(x,z,a^{(K)}(t^{(K)})))]$ to both sides, dividing by $h^{(K)}$ and taking the limit $K\to\infty$ yields
${\mathcal{H}}(x,z,\Delta_{x}\tilde{v}^{(\infty)}(t,\cdot,z),\tilde{\alpha}^{(\infty)}(t,x,z))\leq{\mathcal{H}}(x,z,\Delta_{x}\tilde{v}^{(\infty)}(t,\cdot,z),a)\quad(x,z,a)\in\mathscr{X}_{N}\times[\sigma^{2},\infty)^{d-1}.$
Step 5. Steps 1. to 4. show that given any subsequence $(K_{n})_{n\geq 1}$, we can find a further subsequence $(K_{n_{p}})_{p\geq 1}$ such that $\{v^{(K_{n_{p}})}(\cdot,x,z):(x,z)\in\mathscr{X}_{N}\}$ and $\{\mathfrak{a}^{(K_{n_{p}})}_{y}(\cdot,x,z)/h^{(K_{n_{p}})}:(x,z)\in\mathscr{X}_{N}\}$ converge uniformly to some continuous functions $\tilde{v}^{(\infty)}=\{\tilde{v}^{(\infty)}(\cdot,x,z):(x,z)\in\mathscr{X}_{N}\}$ and $\tilde{\alpha}^{(\infty)}=\{\tilde{\alpha}^{(\infty)}(\cdot,x,z):(x,z)\in\mathscr{X}_{N}\}$, respectively, such that $(\tilde{v}^{(\infty)},\tilde{a}^{(\infty)})$ satisfy the continuous-time NLL equation (2.9). Since the latter has a unique classical solution $(v_{\mathrm{cts}},\alpha_{\mathrm{cts}})$, this proves that the entire sequence $(v^{(K)},\mathfrak{a}^{(K)}/h^{(K)})$ converges uniformly to $(v_{\mathrm{cts}},\alpha_{\mathrm{cts}})$. ∎

### v2

@@ 7 Continuous-time limit
For each $K\geq 1$, set $h^{(K)}:=T/K$ and define the discretization $0=s_{0}^{(K)}<s_{1}^{(K)}<\cdots<s_{k}^{(K)}=T$ of the time interval $[0,T]$ by $s^{(K)}_{k}:=kT/K$. For each $K\geq 1$, let $(v^{(K)},\alpha^{(K)})$ solve ($N$-NLL), which is uniquely defined for sufficiently large $K$, say $K\geq K_{0}$. We extend these functions to $0,T]\times\mathscr{X}_{N}$ by linear interpolation in the time variable and set
$\mathfrak{a}^{(K)}(t,x,z):=\alpha^{(K)}(t,x,z)-e_{x}\in\mathbb{R}^{d},\quad(t,x,z)\in[0,T]\times\mathscr{X}_{N}.$
In the continuous-time problem, let $v_{\mathrm{cts}}:[0,T]\times\mathscr{X}_{N}\to[0,\infty)$ be the unique solution to (2.10) and let $\alpha_{\mathrm{cts}}(t,x,z)=\hat{\alpha}(x,z,\Delta_{x}v_{\mathrm{cts}}(t,\cdot,z))$.
Step 1. We claim that $(v^{(K)},\mathfrak{a}^{(K)}/h^{(K)})_{K\geq K_{0}}$ is a uniformly bounded, equicontinuous family of functions, seen as functions from $0,T]\times\mathscr{X}_{N}$ to $0,\infty)\times\mathbb{R}^{d}$.
Uniform boundedness. By Assumption 2.6, $v^{(K)}$ are uniformly bounded, say, by some $c_{0}<\infty$. By definition, for $0\leq k<K$ and $(x,z)\in\mathscr{X}_{N}$,
$\alpha^{(K)}(s^{(K)}_{k},x,z)=\underset{a\in A(h^{(K)},x)}{\mathrm{arg\,min}}\ \Big\{\ell(x,z,a/h^{(K)})h^{(K)}+\sum_{y\in S(x)}a(y)\mathbb{E}[v^{(K)}(s^{(K)}_{k+1},y,Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k})))$$\Big\}$.
Further, define
$\eta^{(K)}(x,z):=\underset{a\in A(h^{(K)},x)}{\mathrm{arg\,min}}H(x,z,\vec{0},a)=\underset{a\in A(h^{(K)},x)}{\mathrm{arg\,min}}\ell(x,z,a/h^{(K)}),\quad(x,z)\in\mathscr{X}_{N}.$
By Lemma 3.3,
$\frac{1}{h^{(K)}}|\alpha^{(K)}(s^{(K)}_{k},x,z)-\eta^{(K)}(x,z)|_{1}\leq\frac{1}{2\gamma}\,\sum_{y\in S(x)}\mathbb{E}|v^{(K)}(s^{(K)}_{k+1},y,Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k})))|\leq\frac{mc_{0}}{2\gamma}.$
By the definition of $\eta^{(K)}(x,z)$ as the minimizer of $H(x,z,\vec{0},\cdot)$ over $A(h^{(K)},x)$, $\ell(x,z,\eta^{(K)}(x,z)/h^{(K)})\leq\ell(x,z,\bar{a}(x)),$ for $\bar{a}(x)\in\mathbb{R}^{d}$ which can be chosen by $a(x)_{y}=0$ for $y\notin S(x)\cup\{x\}$ and $a(x)_{y}=\sigma^{2}$ for $y\in S(x)\setminus\{x\}$. Note that by Assumption 2.6, the running cost does not depend on $a(x)_{x}$. Now observe that $\ell(x,z,\cdot)$ is a strongly convex function on $\mathbb{R}^{{\mathcal{X}}\setminus\{x\}}\cong\mathbb{R}^{d-1}$ and hence coercive. This implies boundedness of $(|\eta^{(K)}_{-x}(x,z)|_{1}/h^{(K)})_{K\geq 1}$. Using $|a_{-x}-e_{x}|_{1}=2|a_{-x}|_{1}$ for $a\in\Sigma^{d-1}$, we see that $(|\eta^{(K)}(x,z)-e_{x}|_{1}/h^{(K)})_{K\geq 1}$ is bounded by some $\tilde{c}_{0}<\infty$. Putting these two estimates together, we arrive at
$\sup_{K\geq K_{0}}\max_{0\leq k<K}\max_{(x,z)\in\mathscr{X}_{N}}\frac{1}{h^{(K)}}|\alpha^{(K)}(s^{(K)}_{k},x,z)-e_{x}|_{1}\leq\frac{mc_{0}}{2\gamma}+\tilde{c}_{0}=:c_{1}<\infty.$
Note that $(|\alpha^{(K)}(t,x,z)|_{1}/h^{(K)})_{K\geq K_{0}}$ is not a bounded sequence.
Equicontinuity. Let us first observe that, by the same reasoning as in Lemma 7.1, the uniform boundedness of $(\mathfrak{a}^{(K)}/h^{(K)})_{K\geq K_{0}}$ implies that
@@
$|v^{(K)}(s^{(K)}_{k+1},x,z)-v^{(K)}(s^{(K)}_{k},x,z)|$
$\quad\leq|\mathbb{E}^{\alpha^{(K)}\otimes\alpha^{(K)}}[v^{(K)}(s^{(K)}_{k+1},X_{s^{(K)}_{k+1}},Z_{s^{(K)}_{k+1}})]-v^{(K)}(s^{(K)}_{k},x,z)|$
$\quad+|\mathbb{E}^{\alpha^{(K)}\otimes\alpha^{(K)}}[v^{(K)}(s^{(K)}_{k+1},X_{s^{(K)}_{k+1}},Z_{s^{(K)}_{k+1}})]-v^{(K)}(s^{(K)}_{k+1},x,z)|$
$\quad=|\mathbb{E}^{\alpha^{(K)}\otimes\alpha^{(K)}}[v^{(K)}(s^{(K)}_{k+1},X_{s^{(K)}_{k+1}},Z_{s^{(K)}_{k+1}})]-v^{(K)}(s^{(K)}_{k},x,z)|$
$\quad+\!\!\!\!\!\sum_{(x^{\prime},z^{\prime})\neq(x,z)}\!\!\!\!\!\!\!\mathbb{P}^{\alpha^{(K)}\otimes\alpha^{(K)}}[(X_{s^{(K)}_{k+1}},Z_{s^{(K)}_{k+1}})=(x^{\prime},z^{\prime})\,|\,(X_{s^{(K)}_{k}},Z_{s^{(K)}_{k}})=(x,z)]\,|v^{(K)}(s^{(K)}_{k+1},x^{\prime},z^{\prime})-v^{(K)}(s^{(K)}_{k+1},x,z)|$
$\quad\leq|\ell(x,z,\alpha^{(K)}(s_{k}^{(K)},x,z)/h^{(K)})|h^{(K)}+2c_{0}c_{2}\,h^{(K)}\leq(\sup_{D}|\ell(x,z,\cdot)|+2c_{0}c_{2})\,h^{(K)}=:c_{3}h^{(K)}.$
where $D:=\{(a_{y})_{y\in{\mathcal{X}}}\in\mathbb{R}^{d}:a_{y}\in[0,c_{1}/2]\ \text{for all}\ y\neq x\}$. This implies equicontinuity of $(v^{(K)})_{K\geq K_{0}}$. We turn to show the equicontinuity of $(\mathfrak{a}^{(K)})_{K\geq K_{0}}$. As in the proof in Proposition 3.4, by Lemma 3.3,
$|\alpha^{(K)}(s^{(K)}_{k},x,z)$ $\,-\alpha^{(K)}(s^{(K)}_{k+1},x,z)|_{1}$
[…]
$\leq\frac{mL^{(K)}_{k+1}h^{(K)}}{2\gamma}\max_{(x,z)\in\mathscr{X}_{N}}|\alpha^{(K)}_{-x}(s^{(K)}_{k},x,z)-\alpha^{(K)}_{-x}(s^{(K)}_{k+1},x,z)|_{1}+\frac{c_{3}m}{2\gamma}(h^{(K)})^{2},$
@@
$\bigl(2\gamma-mL^{(K)}_{k+1}h^{(K)}\bigr)\max_{(x,z)\in\mathscr{X}_{N}}|\alpha^{(K)}(s^{(K)}_{k},x,z)-\alpha^{(K)}(s^{(K)}_{k+1},x,z)|_{1}\leq c_{3}m(h^{(K)})^{2}.$
@@
[…]
Step 3. For $(x,z)\in\mathscr{X}_{N}$, define
$\varphi^{(K)}_{k}(x,z)$ $:=H(x,z,E(x,z,v^{(K)}(s^{(K)}_{k+1}),\alpha^{(K)}(s^{(K)}_{k})),\mathfrak{a}^{(K)}(s^{(K)}_{k},x,z))-v^{(K)}(s^{(K)}_{k+1},x,z),\quad 0\leq k<K$
$\varphi^{(\infty)}(s,x,z)$ $:={\mathcal{H}}(x,z,\Delta_{x}\tilde{v}^{(\infty)}(s,\cdot,z),\tilde{\alpha}^{(\infty)}(s,x,z))+\mathscr{L}^{\tilde{\alpha}^{(\infty)}(s)}\tilde{v}^{(\infty)}(s)(x,z),\quad s\in[0,T]$
where $\mathscr{L}^{\tilde{\alpha}^{(\infty)}}$ is as in (2.9) with $\beta=\tilde{\alpha}^{(\infty)}$. We claim that
$\lim_{K\to\infty}\ \max_{0\leq k<K}\ \Bigl|\frac{1}{h^{(K)}}\varphi^{(K)}_{k}(x,z)-\varphi^{(\infty)}(s^{(K)}_{k},x,z)\Bigr|=0.$ (7.1)
Indeed, write $\varphi^{(K)}_{k}(x,z)=\ell(x,z,\mathfrak{a}^{(K)}(s^{(K)}_{k},x,z)/h^{(K)})h^{(K)}+\mathrm{(I)}_{k}+\mathrm{(II)}_{k},$ where
$\mathrm{(I)}_{k}$ $:=\sum_{y\in{\mathcal{X}}}\mathfrak{a}^{(K)}_{y}(s^{(K)}_{k},x,z)\,\mathbb{E}[v^{(K)}(s^{(K)}_{k+1},y,Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k})))-v^{(K)}(s^{(K)}_{k+1},x,Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k})))$],
@@
[…]
$\lim_{K\to\infty}\ \max_{0\leq k<K}\ \Bigl|\frac{1}{h^{(K)}}\mathrm{(I)}_{k}-\sum_{y\in{\mathcal{X}}}\tilde{\alpha}^{(\infty)}_{y}(s^{(K)}_{k},x,z)\,(\tilde{v}^{(\infty)}(s^{(K)}_{k},y,z)-\tilde{v}^{(\infty)}(s^{(K)}_{k},x,z))\Bigr|=0.$
@@
$\mathrm{(II)}_{k}=\sum_{z^{\prime}\in\Sigma^{d-1}_{N}}\mathbb{P}[Z^{\prime}(x,z,\alpha^{(K)}(s^{(K)}_{k}))=z^{\prime}]\,\big(v^{(K)}(s^{(K)}_{k+1},x,z^{\prime})-v^{(K)}(s^{(K)}_{k+1},x,z)\big).$
[…]
$\lim_{K\to\infty}\ \max_{0\leq k<K}\ \Bigl|\frac{1}{h^{(K)}}\mathrm{(II)}_{k}-\mathscr{L}^{\tilde{\alpha}^{(\infty)}(s_{k}^{(K)})}\tilde{v}^{(\infty)}(s^{(K)}_{k})(x,z)\Bigr|=0.$
This establishes (7.1).
Step 4. Note that, for every $0\leq k<K$, we have
$\big(v^{(K)}(s^{(K)}_{k+1},x,z)-v^{(K)}(s^{(K)}_{k},x,z)\big)=\varphi^{(K)}_{k}(x,z),\quad(x,z)\in\mathscr{X}_{N}.$ (7.2)
Let $t\in(0,T)$, choose a sequence $t^{(K)}\in\{0,T/K,\ldots,T\}$ converging to $t$ and let $T^{(K)}=t^{(K)}K$. Summing over $0\leq k\leq T^{(K)}-1$ in (7.2), we obtain $(v^{(K)}(t^{(K)},x,z)-v^{(K)}(0,x,z))=\sum_{k=0}^{T^{(K)}-1}\varphi^{(K)}_{k}(x,z)$. Now,
@@
$+\Bigl|\sum_{k=0}^{T^{(K)}-1}\varphi^{(\infty)}(s^{(K)}_{k},x,z)h^{(K)}-\int_{0}^{t}\varphi^{(\infty)}(s,x,z)\,\mathrm{d}s\Bigr|.$
The first term converges to zero by Step 3. while the second term converges to zero by classical arguments. This proves that $\tilde{v}^{(\infty)}(t,x,z)-\tilde{v}^{(\infty)}(0,x,z)=-\int_{0}^{t}\varphi^{(\infty)}(s,x,z)\,\mathrm{d}s$. Hence, $\tilde{v}^{(\infty)}(\cdot,x,z)$ is continuously differentiable on $[0,T]$ with derivative $\varphi^{(\infty)}(\cdot,x,z)$.
Step 5. We claim that
$(\tilde{\alpha}^{(\infty)}(t,x,z))_{y\neq x}=\underset{a\in{\mathcal{R}}(\sigma^{2},x)}{\mathrm{arg\,min}}\ {\mathcal{H}}(x,z,\Delta_{x}\tilde{v}^{(\infty)}(t,\cdot,z),a),\quad(t,x,z)\in[0,T]\times\mathscr{X}_{N}.$
Indeed, for every $0\leq k<K$, $(x,z)\in\mathscr{X}_{N}$ and $a\in A(h^{(K)},x)$, we have
$H(x,z,E(x,z,v^{(K)}(t^{(K)}+h^{(K)}),$ $\,\alpha^{(K)}(t^{(K)})),\mathfrak{a}^{(K)}(t^{(K)},x,z))$
$\leq H(x,z,E(x,z,v^{(K)}(t^{(K)}+h^{(K)}),\alpha^{(K)}(t^{(K)})),a).$
Adding $\mathbb{E}[v(t^{(K)}+h^{(K)},x,Z^{\prime}(x,z,\alpha^{(K)}(t^{(K)})))]$ to both sides, dividing by $h^{(K)}$ and taking the limit $K\to\infty$ yields
${\mathcal{H}}(x,z,\Delta_{x}\tilde{v}^{(\infty)}(t,\cdot,z),\tilde{\alpha}^{(\infty)}(t,x,z))\leq{\mathcal{H}}(x,z,\Delta_{x}\tilde{v}^{(\infty)}(t,\cdot,z),a),\qquad(x,z)\in\mathscr{X}_{N},\,a\in{\mathcal{R}}(\sigma^{2},x).$
[…]